A triangle has a perimeter of 161 miles. Each of the two smaller sides of the triangle is two thirds the length of of the longest side. Find the length of each side of the triangle.
Answer by gonzo(654)

P = 161.
L = larger side.
2/3 * L = smaller side.
P = L + 2*(2/3)*L = L + (4/3 * L) = L * (1 + 4/3) = L * (7/3) = (7/3)*L
(7/3)*L = 161
L = (3/7)*161
L = 69
-----
larger side = 69
smaller side = 2/3 * 69 = 46
-----
69 + 2*(46) = 69 + 92 = 161
-----
sizes prove to be good.
answer is:
-----
larger side = 69
smaller side = 46